if f(x)=3x-2 and g(x)=x^2+1. find f(g(3))
Answer by Nate(3500)

One Way:
f(x) = 3x - 2
f(x) = 3(x^2 + 1) - 2
f(x) = 3x^2 + 1
f(3) = 3*9 + 1 = 28
Another Way:
g(3) = 3^2 + 1 = 10
f(10) = 3*10 - 2 = 28
Answer: 28